You need a 85% alcohol solution. On hand, you have a 90 mL of a 25% alcohol mixture. How much pure alcohol will you need to add to obtain the desired solution?
You will need
mL of pure alcohol
to obtain
mL of the desired 85% solution.

Answer by mananth(16949)

--------percent ---------------- quantity
Alcohol Type I 25 ---------------- 90 ml
Pure alcohol 100 ---------------- x ml
Mixture 85 ---------------- 90 + x ml

25 *90+100x=85( 90+x)
2250+100x=7650+ 85 x
100x-85 x = 7650 - -2250
15 x = 5400
/ 15
x = 360
360 ml of Pure alcohol